# On-Policy Policy Gradient Reinforcement Learning Without On-Policy Sampling

### v2

@@ 7 Discussion
This work has shown that adaptive, off-policy sampling can be used to reduce sampling error in data collected throughout rl training and improve the data efficiency of on-policy policy gradient algorithms. We have introduced an algorithm that scales adaptive off-policy sampling to continuous control rl benchmarks and enables tracking of the changing on-policy distribution. By integrating this data collection procedure into the popular ppo algorithm, the main conclusion of our analysis is that on-policy learning algorithms learn most efficiently with on-policy data, not on-policy sampling. In this section, we discuss limitations of our work and present opportunities for future research.
props builds upon the ros algorithm of Zhong et al. [2022]. While Zhong et al. [2022] focused on theoretical analysis and policy evaluation in small scale domains, we chose to focus on empirical analysis with policy learning in standard rl benchmarks. An important direction for future work would be theoretical analysis of props, in particular whether props also enjoys the same faster convergence rate that was shown for ros relative to on-policy sampling.
A limitation of props is that the update indiscriminately increases the probability of under-sampled actions without considering their importance in gradient computation. For instance, if an under-sampled action has zero advantage, it has no impact on the gradient and need not be sampled. An interesting direction for future work could be to prioritize correcting sampling error for $({\bm{s}},{\bm{a}})$ that have the largest influence on the gradient estimate, i.e., large advantage (positive or negative).
Beyond these more immediate directions, our work opens up other opportunities for future research. A less obvious feature of the props behavior policy update is that it can be used track the empirical data distribution of any desired policy, not only that of the current policy. This feature means props has the potential to be integrated into off-policy rl algorithms and used so that the empirical distribution more closely matches a desired exploration distribution. Thus, props could be used to perform focused exploration without explicitly tracking state and action counts.
@@ 8 Conclusion
In this paper, we ask whether on-policy policy gradient methods are more data efficient using on-policy sampling or on-policy data acquired without on-policy sampling. To answer this question, we introduce an adaptive, off-policy sampling method for on-policy policy gradient learning that collects data such that the empirical distribution of sampled actions closely matches the expected on-policy data distribution at observed states. Our method, Proximal Robust On-policy Sampling (props), periodically updates the data collecting behavior policy so as to increase the probability of sampling actions that are currently under-sampled with respect to the on-policy distribution. Furthermore, rather than discarding collected data after every policy update, props permits more data efficient on-policy learning by using data collection to adjust the distribution of previously collected data to be approximately on-policy. We replace on-policy sampling with props to generate data for the widely-used ppo algorithm and empirically demonstrate that props produces data that more closely matches the expected on-policy distribution and yields more data efficient learning compared to on-policy sampling.
Appendix
[…]
Appendix A Theoretical Results
In this section, we present the proof Proposition 2. We use $d_{m}$, $\pi_{m}$, and $p_{m}$ as the empirical state visitation distribution, empirical policy, and empirical transition probabilities after $m$ state-action pairs have been taken, respectively. That is, $d_{m}(s)$ is the proportion of the $m$ states that are $s$, $\pi_{m}(a|s)$ is the proportion of the time that action $a$ was observed in state $s$, and $p_{m}(s^{\prime}|s,a)$ is the proportion of the time that the state changed to $s^{\prime}$ after action $a$ was taken in state $s$.
@@ Appendix A Theoretical Results
Assume that data is collected with an adaptive behavior policy that always takes the most under-sampled action in each state, $s$, with respect to policy $\pi$, i.e, $a\leftarrow\arg\max_{a^{\prime}}(\pi(a^{\prime}|s)-\pi_{m}(a^{\prime}|s))$. We further assume that ${\mathcal{S}}$ and ${\mathcal{A}}$ are finite. Then we have that the empirical state visitation distribution, $d_{m}$, converges to the state distribution of $\pi$, $d_{\pi}$, with probability $1$:
@@
The proof of this theorem builds upon Lemma 1 and 2 by Zhong et al. [2022]. Note that these lemmas superficially concern the ROS method whereas we are interested in data collection by taking the most under-sampled action at each step. However, as stated in the proof by Zhong et al. [2022], these methods are equivalent under an assumption they make about the step-size parameter of the ROS method. Thus, we can immediately adopt these lemmas for this proof.
Under Lemma 1 of Zhong et al. [2022], we have that $\lim_{m\rightarrow\infty}\pi_{m}(a|s)=\pi(a|s)$ for any state $s$ under this adaptive data collection procedure. We then have the following $\forall s$:
$\lim_{m\rightarrow\infty}d_{m}(s)$ =$\overset{(a)}{=}\lim_{m\rightarrow\infty}\sum_{\tilde{s}}\sum_{\tilde{a}}p_{m}(s|\tilde{s},\tilde{a})\pi_{m}(\tilde{a}|\tilde{s})d_{m}(\tilde{s})$
$=\sum_{\tilde{s}}\sum_{\tilde{a}}\lim_{m\rightarrow\infty}p_{m}(s|\tilde{s},\tilde{a})\pi_{m}(\tilde{a}|\tilde{s})d_{m}(\tilde{s})$
$=\sum_{\tilde{s}}\sum_{\tilde{a}}\lim_{m\rightarrow\infty}p_{m}(s|\tilde{s},\tilde{a})\lim_{m\rightarrow\infty}\pi_{m}(\tilde{a}|\tilde{s})\lim_{m\rightarrow\infty}d_{m}(\tilde{s})$
=$\overset{(b)}{=}\sum_{\tilde{s}}\sum_{\tilde{a}}p(s|\tilde{s},\tilde{a})\pi(\tilde{a}|\tilde{s})\lim_{m\rightarrow\infty}d_{m}(\tilde{s}).$
Here, (a) follows from the fact that the empirical frequency of state $s$ can be obtained by considering all possible transitions that lead to $s$. The last line, (b), holds with probability 1 by the strong law of large numbers and Lemma 2 of Zhong et al. [2022].
@@
Appendix B PROPS Implementation Details
@@
In this appendix, we summarize the behavior of props’s clipping mechanism in Table 1 and additionally describe two relevant implementation details for the props update (Algorithm 2).
@@
Figure 6: In this example, $\pi(\cdot|{\bm{s}})={\mathcal{N}}(0,1)$. After several visits to ${\bm{s}}$, all sampled actions (blue) satisfy $a>0$ so that actions $a<0$ are under-sampled. Without regularization, props will attempt to increase the probabilities of under-sampled action in the tail of target policy distribution (green). The regularization term in the props objective ensures the behavior policy remains close to target policy.
@@
$g({\bm{s}}\,{\bm{a}},{\bm{\phi}},{\bm{\theta}})\in[1-\epsilon_{\textsc{props}}{}{},1+\epsilon_{\textsc{props}}{}{}]$ No $-g({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}})$ $\nabla_{\bm{\phi}}{\mathcal{L}}_{\text{{clip}{}}}$
$g({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}})>1+\epsilon_{\textsc{props}}{}{}$ No $-g({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}})$ $\nabla_{\bm{\phi}}{\mathcal{L}}_{\text{{clip}{}}}$
$g({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}})<1-\epsilon_{\textsc{props}}{}{}$ Yes $-(1-\epsilon_{\textsc{props}}{}{})$ $\mathbf{0}$
@@
Appendix C Computing Sampling Error
We claim that props improves the data efficiency of on-policy learning by reducing sampling error in the agent’s buffer ${\mathcal{D}}$ with respect to the agent’s current (target) policy. To measure sampling error, we use the kl-divergence $D_{\textsc{kl}{}}(\pi_{\mathcal{D}}||\pi_{\bm{\theta}})$ between the empirical policy $\pi_{\mathcal{D}}$ and the target policy $\pi_{\bm{\theta}}$ which is the primary metric Zhong et al. [2022] used to show ros reduces sampling error:
$D_{\textsc{kl}{}}(\pi_{\mathcal{D}}||\pi_{\bm{\theta}})=\mathbb{E}_{{\bm{s}}\sim{\mathcal{D}},{\bm{a}}\sim\pi_{\mathcal{D}}(\cdot|{\bm{s}})}\left[\log\left(\frac{\pi_{\mathcal{D}}({\bm{a}}|{\bm{s}})}{\pi_{\bm{\theta}}({\bm{a}}|{\bm{s}})}\right)\right].$ (7)
We compute a parametric estimate of $\pi_{\mathcal{D}}$ by maximizing the log-likelihood of ${\mathcal{D}}$ over the same policy class used for $\pi_{\bm{\theta}}$. More concretely, we let ${\bm{\theta}}^{\prime}$ be the parameters of neural network with the same architecture as $\pi_{\bm{\theta}}$ train and then compute:
${\bm{\theta}}_{\textsc{MLE}}=\operatorname*{arg\,max}_{{\bm{\theta}}^{\prime}}\sum_{({\bm{s}},{\bm{a}})\in{\mathcal{D}}}\log\pi_{{\bm{\theta}}^{\prime}}({\bm{a}}|{\bm{s}})$ (8)
$D_{\textsc{kl}{}}(\pi_{\mathcal{D}}||\pi_{\bm{\theta}})\approx\sum_{({\bm{s}},{\bm{a}})\in{\mathcal{D}}}\left(\log\pi_{{\bm{\theta}}_{\textsc{MLE}}}({\bm{a}}|{\bm{s}})-\log\pi_{\bm{\theta}}({\bm{a}}|{\bm{s}})\right).$ (9)

### v3

@@ 7 Discussion
This work has shown that adaptive, off-policy sampling can reduce sampling error in data collected throughout RL training and increase the data efficiency of on-policy policy gradient algorithms. In this section, we discuss limitations of our work and present opportunities for future research.
Convergence of PROPS. PROPS builds upon the ROS algorithm of Zhong et al. (2022). While Zhong et al. (2022) focused on theoretical analysis and policy evaluation in low-dimensional domains, we chose to focus on empirical analysis with policy learning in standard RL benchmarks. Proposition 1 and previous convergence results by Zhong et al. (2022) focus on tabular MDPs and provide conceptual grounding for PROPS, but this theory does not immediately extend to continuous MDPs. An important direction for future work would be theoretical analysis of PROPS, in particular whether PROPS also enjoys the same faster convergence rate that was shown for ROS relative to on-policy sampling in continuous MDPs.
Advantage Estimates of Historic Data. While PROPS can reduce sampling error in off-policy historic data, the advantage estimates associated with this data still correspond to historic policies and are thus biased advantage estimates of the current policy. PPO computes advantages using GAE (Schulman et al., 2016) with parameter $\lambda\in[0,1]$ controlling how much the advantage estimates weight multi-step Monte Carlo returns versus value function predictions. In practice, PPO uses a relatively large $\lambda$, typically $\lambda=0.95$, which heavily weights Monte Carlo returns. Since these returns are from trajectories generated by a behavior policy correcting sampling error w.r.t. a historic policy, the returns correspond to the historic policy, not the current one. In other words, PROPS corrects the off-policy bias of the $({\bm{s}},{\bm{a}})$ distribution but not that of the advantage estimates. Empirically, this bias appears minor, as PROPS still improves data efficiency. Nevertheless, it can be mitigated by decreasing $\lambda$ so GAE places greater weight on value function predictions, which are updated to reflect the current policy and will thus produce advantage estimates that better align with the current policy.
Advantage-Weighted Sampling Error Correction. A limitation of PROPS is that the update indiscriminately increases the probability of under-sampled actions without considering their importance in gradient computation. For instance, if an under-sampled action has zero advantage, it has no impact on the gradient and need not be sampled. An interesting direction for future work could be to prioritize correcting sampling error for $({\bm{s}},{\bm{a}})$ that have the largest influence on the gradient estimate, i.e., those with large advantages (positive or negative).
Sampling Error Correction for Off-Policy RL. Beyond these more immediate directions, our work opens up other opportunities for future research. A less obvious feature of the PROPS behavior policy update is that it can be used to track the empirical data distribution of any desired policy, not only that of the current policy. This feature means PROPS has the potential to be integrated into off-policy RL algorithms and used so that the empirical distribution more closely matches a desired exploration distribution. Thus, PROPS could be used to perform focused exploration without explicitly tracking state and action counts.
@@ 8 Conclusion
In this work, we ask if adaptive, off-policy action sampling can produce data that more closely matches the expected on-policy data distribution and improve the data efficiency of on-policy policy gradient methods. To answer this question, we introduce Proximal Robust On-policy Sampling (PROPS), a method that periodically updates a data collection behavior policy to increase the probability of sampling actions that are currently under-sampled with respect to the on-policy distribution. Furthermore, rather than discarding collected data after every policy update, PROPS permits more data-efficient on-policy learning by using data collection to adjust the distribution of previously collected data to be approximately on-policy. We replace on-policy sampling with PROPS to collect data for the popular PPO algorithm and empirically demonstrate that PROPS produces data that more closely matches the expected on-policy distribution and yields more data-efficient learning compared to on-policy sampling.
Acknowledgments
We thank Adam Labiosa, Abhinav Harish, and Brahma Pavse for discussion and comments that improved the final version of this work. This work took place in the Prediction and Action Lab (PAL) at the University of Wisconsin – Madison. PAL research is supported by the National Science Foundation (IIS-2410981) and the Wisconsin Alumni Research Foundation.
Appendix A Core Theoretical Results
In this section, we present the proof of Proposition˜1. We use $d_{m}$, $\pi_{m}$, and $p_{m}$ as the empirical state visitation distribution, empirical policy, and empirical transition probabilities after $m$ state-action pairs have been taken, respectively. That is, $d_{m}(s)$ is the proportion of the $m$ states that are $s$, $\pi_{m}(a|s)$ is the proportion of the time that action $a$ was observed in state $s$, and $p_{m}(s^{\prime}|s,a)$ is the proportion of the time that the state changed to $s^{\prime}$ after action $a$ was taken in state $s$.
@@ Appendix A Core Theoretical Results
Assume that data is collected with an adaptive behavior policy that always takes the most under-sampled action in each state, $s$, with respect to policy $\pi$, i.e., $a\leftarrow\arg\max_{a^{\prime}}(\pi(a^{\prime}|s)-\pi_{m}(a^{\prime}|s))$. We further assume that $\mathcal{S}$} and $\mathcal{A}$} are finite and that the Markov chain induced by $\pi$ is irreducible. Then we have that the empirical state visitation distribution, $d_{m}$, converges to the state distribution of $\pi$, $d_{\pi}$, with probability $1$:
@@
The proof of this theorem builds upon Lemma 1 and 2 by Zhong et al. (2022). Note that these lemmas superficially concern the ROS method whereas we are interested in data collection by taking the most under-sampled action at each step. However, as stated in the proof by Zhong et al. (2022), these methods are equivalent under an assumption that they make about the step-size parameter of the ROS method. Thus, we can immediately adopt these lemmas for this proof.
Under Lemma 1 of Zhong et al. (2022), we have that $\lim_{m\rightarrow\infty}\pi_{m}(a|s)=\pi(a|s)$ for any state $s$ under this adaptive data collection procedure. We then have the following $\forall s$:
[…]
Here, (a) follows from the fact that the empirical frequency of state $s$ can be obtained by considering all possible transitions that lead to $s$. The last line, (b), holds with probability 1 by the strong law of large numbers and Lemma 2 of Zhong et al. (2022).
@@
Appendix B Additional Theoretical Results
In this section, we provide additional theory to describe the relationship between different hyperparameters in PROPS:
  1. 1.
    The amount of sampling error in previously collected data and the size of behavior policy updates.
  2. 2.
    The amount of historic data retained by an agent and the amount of additional data the behavior policy must collect to reduce sampling error.
For simplicity, we first focus on a simple bandit setting and then extend to a tabular RL setting.
Suppose we have already collected $m$ state-action pairs and these have been observed with empirical distribution $\pi_{m}({\bm{a}})$. From what distribution should we sample an additional $k$ state-action pairs so that the empirical distribution over the $m+k$ samples is equal in expectation to $\pi_{\bm{\theta}}$}?
Proposition 3.
Assume that $m$ actions have been collected by running some policy ${\pi_{\bm{\theta}}}({\bm{a}})$ and $\pi_{m}({\bm{a}})$ is the empirical distribution on this dataset. If we collect an additional $k$ state-action pairs using the following distribution, and if $(m+k){\pi_{\bm{\theta}}}({\bm{a}})\geq m\cdot\pi_{m}({\bm{a}})$, then the aggregate empirical distribution over the $m+k$ pairs is equal to ${\pi_{\bm{\theta}}}({\bm{a}})$ in expectation:
$\pi_{b}({\bm{a}})\coloneqq\frac{1}{Z}\left[{\pi_{\bm{\theta}}}({\bm{a}})+\frac{m}{k}\left({\pi_{\bm{\theta}}}({\bm{a}})-\pi_{m}({\bm{a}})\right)\right]$
where $Z=\sum_{{\bm{a}}\in{\mathcal{A}}}\left[{\pi_{\bm{\theta}}}({\bm{a}})+\frac{m}{k}\left({\pi_{\bm{\theta}}}({\bm{a}})-\pi_{m}({\bm{a}})\right)\right]$ is a normalization coefficient.
Proof.
Observe that $(m+k){\pi_{\bm{\theta}}}({\bm{a}})$ is the expected number of times $\bm{a}$} is sampled under $\pi_{\bm{\theta}}$} after $m+k$ steps, $m\cdot\pi_{m}({\bm{a}})$ is the number of times each $\bm{a}$} was sampled thus far, and $k\cdot\pi_{b}({\bm{a}})$ is the expected number of times $\bm{a}$} is sampled under our behavior policy after $k$ steps. We want to choose $\pi_{b}({\bm{a}})$ such that $(m+k){\pi_{\bm{\theta}}}({\bm{a}})=m\cdot\pi_{m}({\bm{a}})+k\cdot\pi_{b}({\bm{a}})$ in expectation.
$\begin{split}(m+k){\pi_{\bm{\theta}}}({\bm{a}})&=k\cdot\pi_{b}({\bm{a}})+m\cdot\pi_{m}({\bm{a}})\\ -k\cdot\pi_{b}({\bm{a}})&=m\cdot\pi_{m}({\bm{a}})-(m+k){\pi_{\bm{\theta}}}({\bm{a}})\\ \pi_{b}({\bm{a}})&=-\frac{m}{k}\pi_{m}({\bm{a}})+\left(\frac{m}{k}+1\right){\pi_{\bm{\theta}}}({\bm{a}})\\ &={\pi_{\bm{\theta}}}({\bm{a}})+\frac{m}{k}\left({\pi_{\bm{\theta}}}({\bm{a}})-\pi_{m}\right({\bm{a}}))\end{split}$
Note that $\pi_{b}({\bm{a}})$ will be a valid probability distribution after normalizing only if
[…]
If $\left(m+k\right){\pi_{\bm{\theta}}}({\bm{a}})<m\cdot\pi_{m}({\bm{a}})$, then prior to collecting additional data with our behavior policy, $\bm{a}$} already appears in our data more times than it would in expectation after $m+k$ steps under $\pi_{\bm{\theta}}$}. In other words, we would need to collect more than $k$ additional samples to achieve zero sampling error (or discard some previously collected samples).
When sampling error is large, behavior policy updates must also be large. Intuitively, the difference ${\pi_{\bm{\theta}}}({\bm{a}})-\pi_{m}({\bm{a}})$ is the mismatch between the true and empirical visitation distributions, so adding this term to $d_{\pi_{\bm{\theta}}}$ adjusts $d_{\pi_{\bm{\theta}}}$ to reduce this mismatch. If ${\pi_{\bm{\theta}}}({\bm{a}})-\pi_{m}({\bm{a}})<0$, then $\bm{a}$} is over-sampled w.r.t. $\pi_{\bm{\theta}}$, and $\pi_{b}$ will decrease the probability of sampling $\bm{a}$}. If ${\pi_{\bm{\theta}}}({\bm{a}})-\pi_{m}({\bm{a}})>0$, then $\bm{a}$} is under-sampled w.r.t. $\pi_{\bm{\theta}}$, and $\pi_{b}$ will increase the probability of sampling $\bm{a}$}. When $|{\pi_{\bm{\theta}}}({\bm{a}})-\pi_{m}({\bm{a}})|$ is small, the optimal $\pi_{b}({\bm{a}})$ requires only a small adjustment from $\pi_{\bm{\theta}}$} (i.e., a small update to the behavior policy is sufficient to reduce sampling error). When $|{\pi_{\bm{\theta}}}({\bm{a}})-\pi_{m}({\bm{a}})|$ is large, the optimal $\pi_{b}({\bm{a}})$ requires a large adjustment from $\pi_{\bm{\theta}}$} (i.e., a large update to the behavior policy is needed to reduce sampling error). We can increase (or decrease) the target KL cutoff $\delta_{\text{PROPS}}$ to allow for larger (or smaller) behavior updates.
When we retain large amounts of historic data, the behavior policy must collect a large amount of additional data to reduce sampling error in the aggregate distribution. The $\frac{m}{k}$ factor implies that how much we adjust $d_{\pi_{\bm{\theta}}}$ depends on how much data we have already collected ($m$) and how much additional data we will collect ($k$). If the $k$ additional samples to collect represent a small fraction of the aggregate $m+k$ samples (i.e. $k<<m$), then $\frac{m}{k}$ is large, and the adjustment to $d_{\pi_{\bm{\theta}}}$ is large. This case generally arises when we retain more and more historic data. If the $k$ additional samples to collect represent a large fraction of the aggregate $m+k$ samples (i.e. $k>>m$), then $\frac{m}{k}$ is small, and the adjustment to $d_{\pi_{\bm{\theta}}}$ is small. This case generally arises when we retain little to no historic data.
The next proposition extends this analysis to the tabular RL setting.
Proposition 4.
Assume that $m$ state-action pairs have been collected by running some policy and $d_{m}({\bm{s}},{\bm{a}})$ is the empirical distribution on this dataset. If we collect an additional $k$ state-action pairs using the following distribution, and if $(m+k)d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})\geq m\cdot d_{m}({\bm{s}},{\bm{a}})$, then the aggregate empirical distribution over the $m+k$ pairs is equal to $d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})$ in expectation:
$d_{b}({\bm{s}},{\bm{a}})\coloneqq\frac{1}{Z}\left[d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})+\frac{m}{k}\left(d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})-d_{m}({\bm{s}},{\bm{a}})\right)\right]$
where $Z=\sum_{({\bm{s}},{\bm{a}})\in{\mathcal{S}}\times{\mathcal{A}}}\left[d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})+\frac{m}{k}\left(d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})-d_{m}({\bm{s}},{\bm{a}})\right)\right]$ is a normalization coefficient.
Proof.
The proof is identical to the proof of Proposition 3, replacing ${\pi_{\bm{\theta}}}({\bm{a}}),\pi_{m}({\bm{a}}),$ and $\pi_{b}({\bm{a}})$ with $d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}}),d_{m}({\bm{s}},{\bm{a}}),$ and $d_{b}({\bm{s}},{\bm{a}})$. ∎
{remark}
[] n practice, we cannot sample directly from the visitation distribution $d_{b}({\bm{s}},{\bm{a}})$ in Proposition 4 and instead approximate sampling from this distribution by sampling from its corresponding policy $\pi_{b}({\bm{a}}|{\bm{s}})=d_{b}({\bm{s}},{\bm{a}})/\sum_{{\bm{a}}^{\prime}\in{\mathcal{A}}}d_{b}({\bm{s}}^{\prime},{\bm{a}}^{\prime})$.
Appendix C PROPS Implementation Details
@@
In this appendix, we describe two relevant implementation details for the PROPS update (Algorithm 2). We additionally summarize the behavior of PROPS’s clipping mechanism in Table 1.
@@
Figure 7: In this example, $\pi(\cdot|{\bm{s}})={\mathcal{N}}(0,1)$. After several visits to $\bm{s}$}, all sampled actions (blue) satisfy $a>0$ so that actions $a<0$ are under-sampled. Without regularization, PROPS will attempt to increase the probabilities of under-sampled actions in the tail of target policy distribution (green). The regularization term in the PROPS objective ensures the behavior policy remains close to target policy.
@@
$g({\bm{s}}\,{\bm{a}},{\bm{\phi}},{\bm{\theta}})\in[1-\epsilon_{\text{PROPS}}{},1+\epsilon_{\text{PROPS}}{}]$ No $g({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}})$ $\nabla_{\bm{\phi}}{\mathcal{L}}_{\text{{clip}{}}}$
$g({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}})>1+\epsilon_{\text{PROPS}}{}$ No $g({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}})$ $\nabla_{\bm{\phi}}{\mathcal{L}}_{\text{{clip}{}}}$
$g({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}})<1-\epsilon_{\text{PROPS}}{}$ Yes $(1-\epsilon_{\text{PROPS}}{})$ $\mathbf{0}$